# CuriGS: Curriculum-Guided Gaussian Splatting for Sparse View Synthesis

### v1

@@ 4.2 Quantitative and Qualitative Results
For large-scale, unbounded scenes from the MipNeRF-360, CuriGS demonstrates robust perceptual quality and strong generalization. Quantitatively, it achieves the best results among the tested methods, with a PSNR of 24.10 dB, SSIM of 0.759, and LPIPS of 0.205 (Tab. 1). The corresponding visual comparisons are shown in Fig. 5. Compared to other baselines, CuriGS exhibits reduced unnatural color shifts and more accurately reconstructs fine details, such as grass and ground textures. These improvements stem from the curriculum strategy, which progressively exposes the model to larger perturbations, while the promotion mechanism filters out inconsistent pseudo-views, ultimately yielding more coherent long-range geometry and perceptually plausible renderings in large-scale scenes.
@@ 4.3 Ablation Study
To better quantify and validate the effectiveness of the proposed method in sparse view reconstruction, we conducted a comprehensive ablation study. Specifically, the experiments focused on two aspects: (1) the performance of the method when the curriculum guidance is removed, and (2) the influence of different numbers of initial training views on reconstruction performance. Since incorporating a depth loss on training views has been widely demonstrated to be beneficial and is not the primary contribution of this study, this component is retained in our baseline model. In the experimental setup, we selected representative samples with typical scene characteristics from each dataset for testing, rather than using all available data.
As reported in Tab. 2, the proposed curriculum guidance leads to consistent improvements in PSNR, SSIM, and LPIPS across all datasets and view configurations.
On the LLFF dataset, curriculum guidance consistently improves performance under all sparse-view conditions. When trained with only two input views, the full strategy achieved a PSNR of 20.94 dB compared to 18.30 dB without guidance, alongside notable increases in SSIM (0.768 vs. 0.601) and decreases in LPIPS (0.194 vs. 0.231). As the number of views increased, overall performance improved, but the model without curriculum guidance still lagged behind.
@@
Table 2: Ablation study on the effect of curriculum guidance under different view numbers across the LLFF, MipNeRF-360, and DTU datasets.
@@ 5 Conclusion
[…]
Supplementary Material
6 Implementation Details
6.1 Dataset Setup
LLFF and MipNeRF-360.
Following standard protocols, we use the LLFF and MipNeRF-360 datasets for evaluating forward-facing and unbounded real-world scenes, respectively. For each dataset, every 8th image is selected as part of the test set, and the remaining images are used for training. To simulate sparse-view conditions, we uniformly subsample a limited number of training views from the available set. All images are downsampled by a factor of 8 for computational efficiency and consistency with previous baselines [zhu2024fsgs, zheng2025nexusgs, park2025dropgaussian, li2024dngaussian, zhang2024cor].
DTU.
Following prior work, we select 15 representative scenes (IDs 8, 21, 30, 31, 34, 38, 40, 41, 45, 55, 63, 82, 103, 110, and 114) from the DTU dataset. For each scene, views 25, 22, and 28 are used as the 3-view training set, while the remaining 25 views are designated for testing. All images are downsampled by a factor of 4, and binary masks are applied to remove background regions, ensuring that both training and evaluation focus solely on the target objects.
6.2 Training Details
LLFF.
For the LLFF scenes, the model is trained for 30,000 iterations. The curriculum-based student sampling phase starts at iteration 3,000 and ends at iteration 24,000. For each teacher view, we generate pseudo-student views using perturbation levels $\sigma\in\{1,2,3,4,5,6,7,8,9,10\}$, with 5 students per level. The quality threshold is set to 0.4.
MipNeRF-360.
In the MipNeRF-360, for the bicycle and stump scenes, the curriculum phase spans from iteration 10,000 to 30,000, with a total of 33,000 iterations. For the remaining scenes, training runs for 30,000 iterations, with the curriculum activation scheduled between iterations 7,000 and 27,000. Student views are generated using perturbation levels $\sigma\in\{2,4,6,8,10\}$, with 3 students per level. The quality threshold is set to 0.45.
DTU.
For the DTU dataset, the model is trained for 13,000 iterations, and curriculum occurs between iterations 2,000 and 12,000. We generate student views using perturbation levels $\sigma\in\{1,2,3,4,5\}$, with 10 students per level. The quality threshold is set to 0.55 for scans 114, 110, and 103, and 0.45 for the remaining scans.
Similar to teacher views, student views are also processed with background masking to exclude irrelevant regions. To obtain accurate masks for pseudo views, we first identify background pixels in each corresponding teacher image. Let $\mathcal{B}_{t}$ denotes the set of background pixel coordinates in the teacher image, and let $\mathbf{I}_{s}(p)$ represents the RGB color of pixel $p$ in the student view. For pixels at corresponding positions in the student view, we compute the color mean $\boldsymbol{\mu}_{s}$ and variance $\boldsymbol{\sigma}_{s}^{2}$ over $\mathcal{B}_{t}$. A student pixel $p$ is classified as background if its color lies within a fixed deviation range around the teacher’s background statistics:
$M_{s}(p)=\begin{cases}1,&\text{if }\left\lVert\mathbf{I}_{s}(p)-\boldsymbol{\mu}_{s}\right\rVert<\tau\,\boldsymbol{\sigma}_{s},\\ 0,&\text{otherwise},\end{cases}$ (9)
where $M_{s}(p)$ denotes the binary background mask for the student view, and $\tau$ is a constant threshold controlling the tolerance range. This adaptive masking strategy enables robust background suppression and maintains mask consistency despite geometric perturbations between teacher and student cameras.
7 Extended Ablation Analysis
We previously demonstrated that the introduction of student views effectively alleviates overfitting in sparse-view reconstruction. To further quantify this effect, we conducted a detailed training analysis on a representative DTU scene, continuously monitoring the evolution of PSNR, SSIM, and LPIPS throughout the optimization process for both the full CuriGS framework and an ablated version without curriculum guidance, as shown in Fig. 7. As training progresses, the model without curriculum guidance exhibits clear signs of overfitting—its performance metrics gradually deteriorate as it memorizes limited viewpoints rather than improving scene generalization. In contrast, the full CuriGS strategy maintains steady performance gains across all metrics, particularly in the later training stages, demonstrating that the curriculum-based student-view integration provides sustained and stable supervision that effectively prevents overfitting and promotes consistent convergence.
8 Discussion
Figure 8: Visualization of a teacher view and its associated student-view group generated with an example perturbation magnitude. From left to right, the rendered results of the students are shown at different training stages, demonstrating how visual fidelity and structural coherence improve over time. Scores are computed as a combination of multiple metric losses, with lower values indicating better quality The best student is labeled with a blue box.
The motivation behind our curriculum-based learning strategy is to effectively expand the sparse set of training views. In sparse-view scenarios, limited camera coverage restricts supervision diversity, making it crucial to introduce additional, yet reliable, viewpoints. Therefore, a key challenge lies in how to efficiently select and utilize high-quality student views. To this end, we design a progressive training mechanism where student views with different perturbation magnitudes are gradually introduced to assist optimization, and the best-performing students are continuously tracked for potential inclusion in the training set. To evaluate student quality, we integrate three complementary metrics—LPIPS, SSIM, and a no-reference image-quality score—forming a unified criterion that compensates for the absence of explicit ground-truth supervision in pseudo views. LPIPS and SSIM capture perceptual and structural similarity with the corresponding teacher render, ensuring semantic and visual consistency, while the no-reference score enforces overall image plausibility. Empirically, this composite evaluation yields a stable and balanced metric for identifying reliable student views, effectively bridging the supervision gap between real and synthetic viewpoints.
When training reaches a perturbation-switching stage, the best student from each group undergoes a final quality evaluation. If its rendered result surpasses a predefined threshold, the view is promoted into the training set. At this stage, only the image-quality score is used for evaluation, since LPIPS and SSIM tend to favor students closer to the teacher view and thus provide limited diversity. Allowing moderate geometric or appearance deviations encourages the model to explore broader scene variations, ultimately improving generalization and robustness. Fig. 8 visualizes a teacher view and its corresponding student group at a specific perturbation level, showing progressive refinement of student renders and the selection of the best-performing candidate.
9 Additional Visualization Results
To further illustrate the effectiveness and robustness of our approach, we provide additional qualitative comparisons on the LLFF, Mip-NeRF360, and DTU datasets. As shown in Figs. 9, 10 and 11, CuriGS produces visually coherent and geometrically stable reconstructions across diverse scene types, ranging from real forward-facing captures to unbounded outdoor environments and object-centric setups. These results collectively demonstrate the versatility and generalization ability of CuriGS across different domains and levels of sparsity.
Figure 9: Per-scene visualization results on LLFF.
Figure 10: Per-scene visualization results on MipNeRF-360.
Figure 11: Per-scene visualization results on DTU.

### v2

@@ 4.4 Quantitative and Qualitative Results
For large-scale, unbounded scenes from the MipNeRF-360, CuriGS demonstrates robust perceptual quality and strong generalization. Quantitatively, it achieves the best results among the tested methods, with a PSNR of 24.21 dB, SSIM of 0.761, and LPIPS of 0.202 (Tab. 1). The corresponding visual comparisons are shown in Fig. 5. Compared to other baselines, CuriGS exhibits reduced unnatural color shifts and more accurately reconstructs fine details, such as grass and ground textures. These improvements stem from the curriculum strategy, which progressively exposes the model to larger perturbations, while the promotion mechanism filters out inconsistent pseudo-views, ultimately yielding more coherent long-range geometry and perceptually plausible renderings in large-scale scenes.
@@ 4.5 Ablation Study
To better quantify and validate the effectiveness of the proposed CuriGS framework in sparse-view reconstruction, we conduct a comprehensive ablation study.
Effectiveness of Curriculum Guidance.
The overall impact of curriculum guidance across varying datasets and view numbers is demonstrated through consistent and significant improvements in PSNR, SSIM, and LPIPS under diverse sparse-view configurations, as reported in Tab. 2. On the LLFF dataset, curriculum guidance consistently improves performance under all sparse-view conditions. When trained with only two input views, the full strategy achieved a PSNR of 20.94 dB compared to 18.30 dB without guidance, alongside notable increases in SSIM (0.768 vs. 0.601) and decreases in LPIPS (0.194 vs. 0.231). As the number of views increased, overall performance improved, but the model without curriculum guidance still lagged behind.
@@
Table 2: Ablation study on the effect of curriculum guidance.
@@
Effectiveness of Loss Components.
To specifically address the structural integrity of our optimization process, we ablate the core loss components formulated in our method: the dynamic reconstruction loss, the anchor loss, and the student-view regularization. First, removing the anchor loss on original views leads to noticeable geometric drift and a drop in overall PSNR, demonstrating its critical role in anchoring the foundational 3D structure against perturbations. Second, excluding the student-view regularization results in the model severely overfitting to the sparse inputs, as it fails to penalize inconsistent geometric priors introduced by unconstrained pseudo-views. Finally, the full triple-loss structure ensures that the dynamic reconstruction leverages the progressive perturbations effectively while maintaining strict fidelity to the ground-truth observations, yielding the most photorealistic and geometrically stable novel-view synthesis.
Table 3: Ablation study on the contribution of each loss component across the LLFF, MipNeRF-360, and DTU datasets.
[…]
LLFF 3 / Full 22.51 0.834 0.150
3 / w/o Anchor Loss 21.84 0.795 0.178
3 / w/o Student Reg. 21.48 0.782 0.190
M-360 16 / Full 21.23 0.781 0.189
16 / w/o Anchor Loss 20.88 0.768 0.194
16 / w/o Student Reg. 20.65 0.760 0.198
DTU 3 / Full 22.65 0.947 0.050
3 / w/o Anchor Loss 21.97 0.935 0.058
3 / w/o Student Reg. 20.82 0.929 0.061
Figure 7: Visualization of a teacher view and its associated student-view group under a specific perturbation magnitude. From left to right, the rendered results of the students are shown across different training stages, illustrating the progressive enhancement of visual fidelity and structural coherence. The composite evaluation scores (where lower indicates better quality) are utilized to track performance, with the optimal student candidate highlighted by a blue box.
4.6 Analysis of Pseudo-View Selection Mechanism
In extreme sparse-view scenarios, the fundamental bottleneck lies in the severe lack of multi-view supervision. While expanding the training set via curriculum-guided pseudo-views can mitigate this issue, the unchecked inclusion of poorly rendered views inevitably degrades the underlying geometric fidelity. Therefore, a critical challenge is establishing a robust and dynamic criterion to efficiently evaluate and select high-quality student candidates.
To bridge the supervision gap caused by the absence of explicit ground truth for pseudo-views, we formulate a composite evaluation metric during the continuous tracking phase. This metric elegantly integrates full-reference metrics (LPIPS and SSIM) with a no-reference image-quality score. Specifically, LPIPS and SSIM measure the structural and perceptual deviations from the corresponding teacher view, ensuring that the student maintains strict semantic consistency with the known observations. Simultaneously, the no-reference score penalizes synthetic rendering artifacts, guaranteeing overall image plausibility. Empirically, this tripartite criterion serves as a stable proxy for supervision, effectively screening out degraded viewpoints during the early optimization stages.
Crucially, our strategy adaptively shifts when the training reaches a perturbation-switching stage. During the final evaluation to promote a student into the permanent training pool, we deliberately decouple the criteria by relying solely on the no-reference quality score. The rationale behind this design is that heavily weighting LPIPS and SSIM at the promotion stage inherently biases the selection toward students spatially closest to the teacher, thereby stifling viewpoint diversity. By tolerating moderate geometric or appearance deviations—provided the standalone rendering quality remains high—the model is compelled to explore broader scene variations. As visualized in Fig. 7, this dynamic selection mechanism not only progressively refines the student renders but also successfully identifies the optimal candidate that balances structural coherence with informative viewpoint expansion, ultimately enhancing the generalization and robustness of the 3D representation.
4.7 Analysis of Training Dynamics
To further investigate the dynamic impact of our curriculum-guided expansion, we monitor the evolution of PSNR, SSIM, and LPIPS metrics throughout the training process on a representative DTU scene. As depicted in Fig. 8, the baseline without curriculum guidance exhibits a distinctive performance decay in the later stages of optimization—a hallmark of overfitting where the model memorizes sparse training viewpoints at the expense of general scene structure. In contrast, CuriGS maintains a steady and upward trajectory across all evaluation metrics. This sustained improvement demonstrates that the progressive integration of validated student views provides a continuous and reliable supervisory signal, effectively anchoring the optimization process and ensuring consistent convergence toward a high-fidelity representation.